Show that R^n and C^n are not compact.


1
Expert's answer
2021-07-05T16:51:47-0400

First of all let us show that "\mathbb{R}^n" is not compact. There are two possible definitions of compactness for metric spaces (Bolzano-Weierstrass sequence property and Borel-Lebesgue open cover property), let us disprove both of them for "\mathbb{R}^n".

  • Let us consider a sequence "u_n=\begin{pmatrix} n \\0 \\ ... \\ 0 \end{pmatrix}" for "n\in \mathbb{N}". This sequence does not admit a converging subsequence, as for any subsequence the first coordinate diverges to infinity. Therefore, "\mathbb{R}^n" does not admit the Bolzano-Weierstrass property and thus is not compact (as it is metric).
  • Let us consider an open cover "U_n=B(0,n)" by the open balls centered at origin with a growing radius. This open cover does not admit a finite subcover, as for any finite subcover we have "\bigcup_{n\leq N} U_n \subseteq B(0,N+1)\neq \mathbb{R}^n". Therefore, "\mathbb{R}^n" does not admit the Borel-Lebesgue property and thus is not compact.

To prove that "\mathbb{C}^n" is not compact, we can either use the former approach or use a simpler argument : if "\mathbb{C}^n" were compact, then "\mathbb{R}^n\subseteq \mathbb{C}^n" being a closed subset of a compact space, would be compact too. However we have just proven the contrary, and thus "\mathbb{C}^n" is not compact.